A firm uses straight line depreciation for fixed assets with an estimated useful life of 12 years for its financial statements and 8 years for taxable income.

●    Equipment is bought for 500 on December 31, 20X0.
●    The firm’s corporate income tax payments are 65 in 20X1, 69 in 20X2, and 57 in 20X3.

●    In 20X1 and 20X2, the corporate tax rate is expected to be 19% for all years.
●    On January 1, 20X3, legislation is enacted that reduces the tax rate to 10% for 20X4 and later years.


1. Depreciation on the equipment causes the only deferred tax assets or liabilities for the firm.

For the year 20X1

  1. What is the carrying value of the equipment on December 31, 20X1?

    b. What is the tax basis of the equipment on December 31, 20X1?

    c.What is the deferred tax asset (liability) at December 31, 20X1?

    d. What is the tax expense in 20X1?

    e. Carrying value for the Equipment Dec 20x2 will be ?
Answer #1

a) Carrying value of equipment

= Orginal cost-Accumulated depreciation

= 500- (W.N.1)41.67

= 458.33

W.N.1

= Cost of asset/no. of life

= 500/12

= 41.67

b) tax basis of equipment

= cost of asset*corporate tax rate

= 500*19%

= 95

c) Deferred tax asset (Liability)

An item recorded on the financial statements denotes the excess payment of the taxes in accordance with the equipment is called a deferred tax asset.

Deferred tax asset

= taxes paid - actual taxe paid

= 65-95

= -30 is a liability

d) tax expenses

Estimated depreciation expense*corporate tax rate

= 41.67*19%

= 7.91

e) Carrying value of equipment

= Orginal cost-Accumulated depreciation

= 500- (W.N.1)41.67

= 458.33

W.N.1

= Cost of asset/no. of life

= 500/12

= 41.67